Two of the most expensive movies ever made were Titanic and Water world. The cost to make Titanic exceeded the cost to make Water world by $25 million. The combined cost to make the two movies was $375 million. Find the cost of making each of these movies. Thanks im having trouble figuring this out.
Found 2 solutions by edjones, stanbon:
Answer by edjones(8007)

let x=cost of making water world m=million
x+x+25m=375m
2x+25m=375m
2x=350m
x=$175,000,000 cost of making water world
x+25m=$200,000,000 cost of making titanic
Ed

Answer by stanbon(75887)

The cost to make Titanic exceeded the cost to make Water world by $25 million. The combined cost to make the two movies was $375 million. Find the cost of making each of these movies.
------------------
Let the Titanic cost be "x" million dollars:
Water World cost must be "x+25" million dollars:
----------------
EQUATION:
x + x+25 = 375
2x = 350
x = 175 million dollars (cost of Titanic)
x+25 = 200 million dollars (cost of Water World)
=======================
Cheers,
Stan H.